OpenStudy (anonymous):

Write the point-slope form of an equation for the line that passes through the point (5,-3) with the given slope m = ½

OpenStudy (lucaz):

the point slope form is \[\large{y-y_1=m(x-x_1)}\] you have the slope the the point (5,-3), replace them in the point-slope form

OpenStudy (anonymous):

i really dont get it still

OpenStudy (lucaz):

from the point (5,-3) 5 is x and -3 is y, replace x1 by 5 and y1 by -3, and m by the slope 1/2

OpenStudy (anonymous):

would my awnser be y= 1/2x +5.5 or y-3=1/2(x-5)

OpenStudy (anonymous):

yes

OpenStudy (britanyashton):

m=slope the slope is like how dense a line is on a four column graph. y with a small 1 is the first number of the (x, y) sequence you have http://www.teachertube.com/search_result.php?query=point+slope+form&type=all&safe=high&search-type=all&head_curriculum=0&head_content_area=0&head_subject=0&head_category=0&head_skill=0 watch one of these videos a couple times. maybe once a day for however long you need to til you got it down. its usually easier to watch a video explaining than looking at a book.

OpenStudy (lucaz):

answer would be y-(-3)=1/2(x-(5)) y+3=1/2x-5/2 or y=1/2x-9/2 or y=1/2x-4.5